OpenStudy (dtan5457):

Which equation has a diameter with endpoints (3,-2) and (-5,8)?

OpenStudy (dtan5457):

1. (x+1)^2+(y-3)^2=41 2.(x-1)^2+(y+3)^2=41 3. (x+3)^2+(y-2)^2=164 4.(x-3)^2+(y+2)^2=2 (square root of 41)

OpenStudy (anonymous):

(x-x1)(x-x2)+(y-y1)(y-y2)=0

OpenStudy (dtan5457):

1. (x+1)^2+(y-3)^2=41?

OpenStudy (anonymous):

(x-3)(x+5)+(y+2)(y-8)=0

OpenStudy (anonymous):

1.siplify it.

OpenStudy (dtan5457):

my answer is wrong?

OpenStudy (anonymous):

or find the middle points of the extremities of diameter,it is center.

OpenStudy (dtan5457):

yes i found midpoint which is -1,3 and i automatically picked 1.

OpenStudy (anonymous):

your answer is correct

OpenStudy (anonymous):

i was telling you other method. second method is easy.

OpenStudy (danjs):

(x - h)^2 + (y - k)^2 = r^2 to find the radius r, compute the distance between those two points and take half of that. To get the center (h,k) Use the midpoint of that diameter.